# The Open mulTiwavelength Transient Event Repository (OTTER): Infrastructure Release and Tidal Disruption Event Catalog

### v1

@@ Abstract
Multiwavelength analyses of astrophysical transients are essential for understanding the physics of these events. To make such analyses more efficient and effective, we present the Open mulTiwavelength Transient Event Repository (OTTER), a publicly available catalog of published transient event metadata and photometry. Unlike previous efforts, our data schema is optimized for the storage of multiwavelength photometric datasets spanning the entire electromagnetic spectrum. Open source software, including an application programming interface (API) and web application, are available for viewing, accessing, and analyzing the dataset. For the initial release of OTTER, we present the largest ever photometric archive of tidal disruption events (TDEs), including $\gtrsim 80,000$ observations of 232 TDEs spanning from radio to X-rays wavelengths. We demonstrate the power of this infrastructure through four example analyses of the TDE population. We plan to maintain this dataset as more TDEs are discovered in the future and encourage other users to contribute by uploading newly published data via our web application. The infrastructure was built with the goal of archiving additional transient data (supernovae, gamma-ray bursts, fast blue optical transients, fast radio bursts, etc.) in the future. The web application is available at https://otter.idies.jhu.edu and the API documentation is available at https://astro-otter.readthedocs.io.
@@ 1 Introduction
The study of astronomical transients presents a unique view of the Universe through an indispensable high-energy laboratory that is impossible to recreate on Earth. Examples of transient events include 1) core collapse supernovae (CCSNe), when a massive star ends its life in a violent explosion (radiated isotropic energy, $E_{\rm iso,\,rad}\sim 10^{51}$ erg; Colgate & White, 1966; Woosley & Weaver, 1995; Smartt, 2009); 2) thermonuclear supernovae (e.g., Type Ia SNe), when a degenerate white dwarf star explodes ($E_{\rm iso,\,rad}\sim 10^{51}$ erg; Colgate & White, 1966; Maoz et al., 2014); 3) gamma-ray bursts (GRBs; Kouveliotou et al., 1993; Norris et al., 1984), hypothesized to occur when a massive star collapses and produces a jet (Heger et al., 2003) and/or when two neutron stars merge (Eichler et al., 1989; Narayan et al., 1992; Abbott et al., 2017; Savchenko et al., 2017; Goldstein et al., 2017; Fong et al., 2017, $E_{\rm iso,\,rad}\sim 10^{54}$ erg); and 4) tidal disruption events (TDEs; $E_{\rm iso,\,rad}\sim 10^{50-54}$ erg), when a star is torn apart by a supermassive black hole (SMBH) (Hills, 1975; Rees, 1988).
@@
OTTER is a successor to the original Open Astronomy Catalogs (OAC; i.e. the Open Supernovae Catalog and Open TDE Catalog; Guillochon et al., 2017; Auchettl et al., 2017), which have not been maintained since $\sim 2022$. Building off the OACs, we designed our infrastructure to both inherently store multiwavelength data and scale efficiently with large all sky surveys. In particular, we have built a flexible dataset structure (i.e., a “data schema”) to explicitly store information about the nuances related to X-ray and radio data in the catalog (e.g., the fact that X-ray fluxes are model dependent, see §2.3.3). Additionally, we incorporate the ability for each transient to have multiple classifications with descriptive flags to allow for appropriate sample selection. This is important to account for transients that change their classifications because of new observational evidence (e.g., AT 2017bcc, which was initially classified as a TDE but later reclassified as an ambiguous nuclear transient; Barbarino et al., 2017; Ridley et al., 2024) or transients that appear to transition between classifications over time (e.g., SN PTF11iqb, which observationally evolved from a Type IIn to Type IIP/L at late times; Smith et al., 2015).
@@
For this first release, we focus on ingesting and analyzing data from tidal disruption events (TDEs). Observations of TDEs exist across the electromagnetic spectrum, making them a good starting point for the multiwavelength capabilities of OTTER. As a result, each transient currently in OTTER has, at least at one time, been classified as a TDE. To illustrate the types of analyses facilitated by OTTER, we also present several worked examples focused on exploring the TDE population and comparing to theory.
This paper is organized as follows. In §2 we describe the dataset schema, the data cleaning process, the database backend, and the classification flagging. In §3, we outline the website and the API, and describe how to contribute new datasets to the catalog as they get published. In §4, we summarize the current state of the dataset. In §5, we outline a few applications of the OTTER dataset and API to the TDE population currently stored in OTTER. We conclude and discuss future directions in §6. Appendix A lists the metadata and references for all of the transients currently in OTTER. Appendix B extends §2 with a detailed description of the dataset schema. Appendix C includes additional plots of the current OTTER dataset.
2 Methodology & Data Organization
@@
    A complete dataset of transient events that have at least one associated refereed publication, including their photometry and meta data (names, coordinates, distances, etc.). In this work we start with TDE data but with the plan to expand OTTER to include other transient events in the future.
@@
    Cleaned Dataset: Similar to the OAC, and for longevity, we store all of our data in a human readable JSON format. The raw JSON files are currently available on GitHub [^2] https://github.com/astro-otter/otter-backups and will be made available on Zenodo as the dataset size grows. These JSON files are described in §2.1. To curate a clean dataset of TDE photometry, we pull raw datasets from papers, public catalogs, and data uploaded by external users to the OTTER website and run them through our data cleaning pipeline, described in §2.3.
@@
    Backend Database: We use a database server to make the files programmatically accessible for both the API and the website. This is described in §2.2. This is hosted by Johns Hopkins’ SciServer [^3] https://www.sciserver.org/.
@@
    Website: The OTTER website provides an easy view of the data available in OTTER. This is described in §3.1. This is also hosted by Johns Hopkins’ SciServer.
@@
    API: There is a Python API, built on the more general RESTful API [^4] https://docs.arangodb.com/stable/develop/http-api/, that allow user access to the cleaned dataset. This is described in §3.2 and in the API documentation [^5] https://astro-otter.readthedocs.io/. The API documentation also includes multiple Python Jupyter notebooks with examples of using the API to query the database [^6] https://astro-otter.readthedocs.io/en/latest/examples/welcome.html.
2.1 Data Schema
@@ 2.2 Data Backend
2.3 Data Sources & Cleaning Pipeline
To create the initial OTTER TDE catalog, we compile data from multiple existing catalogs and individual papers. In addition to cleaning the data into a standard format, each component of each dataset is appropriately cited with an ADS bibcode in the individual JSON files. All of these citations are given in Table LABEL:tab:refs. In §2.3.1 through §2.3.5 we describe dataset specific nuances for the data cleaning process.
@@ 2.3 Data Sources & Cleaning Pipeline
Throughout the data cleaning process, we deduplicate identical data (metadata and photometry) by appending the new citation to the previously stored data. This ensures that all appropriate works are cited when the data is used. An intentional artifact of only deduplicating identical datasets is that it does not remove different reductions of the same photometry, allowing the user to choose which flux value they prefer. To prevent users from accidentally using multiple photometry points produced from different reductions of the same data, we provide a default deduplication algorithm as part of the Python API. See §3.2 for a discussion of this algorithm.
2.3.1 The Open TDE Catalog
The Open TDE Catalog was, like the Open Supernova Catalog, part of the AstroCats movement to create a standardized data format for public time domain data, including published and unpublished datasets (Guillochon et al., 2017; Auchettl et al., 2017). Until $\sim 2022$, this was the most comprehensive catalog of TDE data, with data on $\sim 100$ TDEs. The data organization in OTTER is influenced by the AstroCats movement, which also used the JSON format for their backend data storage.
@@ 2.3.1 The Open TDE Catalog
2.3.2 Ultraviolet, Optical, and Infrared Data (UVOIR)
The so-called “Curated Optical TDE Catalog” is a catalog of optically-selected TDE photometry available at https://github.com/sjoertvv/manyTDE (Mummery et al., 2024b). This dataset includes more recent optical/UV/IR photometry for optically-selected TDEs, most of which were discovered after the deprecation of The Open TDE Catalog. Most of this photometry is from the Zwicky Transient Facility (ZTF; van Velzen et al., 2019) forced photometry server, WISE (Wright et al., 2010; Mainzer et al., 2011), and the Neil Gehrels Swift Ultraviolet/Optical Telescope (Swift UVOT), with some additional photometry from other optical all sky surveys (Shappee et al., 2014; Tonry et al., 2018c; Blagorodnova et al., 2017a; Gezari et al., 2012a). The catalog was initially assembled for the analysis performed in Mummery et al. (2024b), using methods outlined in van Velzen et al. (2019) and van Velzen et al. (2019), but has grown since then. All of the data is host subtracted and we store the appropriate host subtraction in the OTTER schema. Every time our data cleaning pipeline runs it pulls changes from The Curated Optical TDE Catalog and incorporates the photometry into OTTER.
@@ 2.3.2 Ultraviolet, Optical, and Infrared Data (UVOIR)
Masterson et al. (2024) performed an archival search of WISE Infrared (IR) data for flares associated with potential tidal disruption events and found 12 new IR bright TDEs. We include their entire photometric dataset, spanning from IR to X-rays, in OTTER as well as specific metadata information like coordinates, redshift, and best-fit disruption time. This dataset was provided through private communication with the authors of Masterson et al. (2024).
@@ X VLA (8, 12]
2.3.5 The Transient Name Server
@@
2.4 Classification Confidence Flagging
Accurately and efficiently classifying transients, particularly rare transients like TDEs, remains a complex problem and a matter of intense discussion. For example, in the case of TDEs, while the gold standard in the field is often taken to be the existence of a high signal-to-noise optical spectrum with characteristic emission lines (e.g., Arcavi et al. 2014), this excludes TDEs that are discovered at other wavelengths (e.g., X-ray, radio) and do not show any optical emission — even though many of these objects are also widely accepted as TDEs (e.g., Auchettl et al., 2017; Anderson et al., 2020). To help with this ambiguity and facilitate the curation of “reliable” population sub-samples for different science cases, we provide numerous flags for the classifications associated with any given transient. The classification flags described in this subsection are applicable to most transients that will eventually be included in OTTER, including both TDEs and SNe. However, with the addition of some transient events (e.g., GRBs) in the future, other, more relevant, classification flags may become necessary. These flags also prepare us for a future where many transients will not be optically spectroscopically classified, in which case the classification may remain ambiguous indefinitely.
@@ 2.4 Classification Confidence Flagging
Based on this flagging schema, we can outline the automated algorithm for deriving meta-level flags describing how each transient was classified. Figure 2 is a flow chart that shows this algorithm. Currently, we store three such meta-level flags. The first is simply the discovery method of the transient (discovery_method). The discovery_method flag is currently either “xray”, “uvoir”, or “radio”, but it could be extended to also include gamma-rays, neutrinos, or gravitational waves as future discovery methods. This flag can be particularly useful for selecting TDE samples as the discovery wavelength regime is tied to some astronomers’ confidence that the event is a TDE rather than, for example, a SNe.
@@
Figure 3 shows the distribution of the maximum confidence flags associated with each transient event (i.e., TDE or TDE candidate) stored in OTTER. This distribution demonstrates that most ($\sim 150$) the events in OTTER are classified from optical spectroscopy and about a third as many are only photometrically classified. The most pure sample of TDEs in OTTER probably has the flags unambiguous $\rm true$} and spec_classed $\geq 1$. However, selecting on these flags does bias the sample to optically-selected TDEs, even though there are other highly probable TDEs in the catalog. As a result, we leave the exact sample selection to the discretion of the individual user.
@@
3 Data Availability & Contributing Data
As described in §2, two of the primary goals of OTTER are 1) allow for the data to be easily accessed and 2) make it easy to contribute data after it is published, with the hope that others will make this contribution part of their typical publishing workflow. As such, we developed two ways to interface with the OTTER dataset, neither of which require any account registration or login: 1) an open access website for easily viewing and uploading datasets and 2) an API for programmatic access to the data. In the following subsections, we describe this infrastructure.
@@ 3.1 OTTER Website
Each individual transient page includes the names assigned to the transient, a table of its basic properties, and plots of its light curve and spectral energy distribution. A sample of one of these pages is shown in Figure 5. The properties table lists what we have deemed to be the “default” values for the various transient properties that are available. For most properties, if multiple values are reported in the literature, then the default value is determined by the value that has the most citations. There are two exceptions: 1) the default name of the transient, for which we select the International Astronomical Union convention (i.e. TNS or TNS-like) names over other transient aliases used in the literature (which are also listed in the table) and 2) the default classification, which is based on the highest confidence flag161616see §2.4 for a discussion of the confidence flagging and the biases that this default classification algorithm may introduce.. The light curve and spectral energy distribution plots are shown if photometry is publicly available.
3.1.1 Uploading Data via the OTTER Website

### v2

@@ Abstract
Multiwavelength analyses of astrophysical transients are essential for understanding the physics of these events. To make such analyses more efficient and effective, we present the Open mulTiwavelength Transient Event Repository (OTTER), a publicly available catalog of published transient event metadata and photometry. Unlike previous efforts, our data schema is optimized for the storage of multiwavelength photometric datasets spanning the entire electromagnetic spectrum from multiple published sources. Open source software, including an application programming interface (API) and web application, are available for viewing, accessing, and analyzing the dataset. For the initial release of OTTER, we present the largest ever photometric archive of tidal disruption event (TDE) candidates, including $\gtrsim$ observations of 240 TDE candidates spanning from radio to X-ray wavelengths. We demonstrate the power of this infrastructure through four example analyses of the TDE population. We plan to maintain this dataset as more TDE candidates are proposed in the future and encourage other users to contribute by uploading newly published data via our web application. The infrastructure was built with the goal of archiving additional transient data (supernovae, gamma-ray bursts, fast blue optical transients, fast radio bursts, etc.) in the future. The web application is available at https://otter.idies.jhu.edu and the API documentation is available at https://astro-otter.readthedocs.io.
@@ I Introduction
The study of astronomical transients presents a unique view of the Universe through an indispensable high-energy laboratory that is impossible to recreate on Earth. Examples of transient events include 1) core collapse supernovae (CCSNe), when a massive star ends its life in a violent explosion (radiated isotropic energy, $E_{\rm iso,\,rad}\sim 10^{51}$ erg; Colgate & White, 1966; Woosley & Weaver, 1995; Smartt, 2009); 2) thermonuclear supernovae (e.g., Type Ia SNe), when a degenerate white dwarf star explodes ($E_{\rm iso,\,rad}\sim 10^{51}$ erg; Colgate & White, 1966; Maoz et al., 2014); 3) gamma-ray bursts (GRBs; Kouveliotou et al., 1993; Norris et al., 1984), hypothesized to occur when a massive star collapses and produces a jet (Heger et al., 2003) and/or when two neutron stars merge (Eichler et al., 1989; Narayan et al., 1992; Abbott et al., 2017; Savchenko et al., 2017; Goldstein et al., 2017; Fong et al., 2017, $E_{\rm iso,\,rad}\sim 10^{54}$ erg); and 4) tidal disruption events (TDEs; $E_{\rm iso,\,rad}\sim 10^{50-54}$ erg), when a star is torn apart by a supermassive black hole (SMBH; Hills, 1975; Rees, 1988).
@@
OTTER is a successor to the original Open Astronomy Catalogs (OAC; i.e. the Open Supernovae Catalog and Open TDE Catalog; Guillochon et al., 2017; Auchettl et al., 2017), which have not been maintained since $\sim 2018$. Building off the OACs, we designed our infrastructure to both inherently store multiwavelength data and scale efficiently with large all sky surveys. In particular, we have built a flexible dataset structure (i.e., a “data schema”) to explicitly store information about the nuances related to X-ray and radio data in the catalog (e.g., the fact that X-ray fluxes are model dependent, see §II.3.3). Additionally, we incorporate the ability for each transient to have multiple classifications with descriptive flags to allow for ease of sample selection (e.g., a flag that tells the user if the event is spectroscopically classified, or another flag that tells the user if the classification is debated). This is important to account for transients that change their classifications because of new observational evidence (e.g., AT 2017bcc, which was initially classified as a TDE but later reclassified as an ambiguous nuclear transient; Barbarino et al., 2017; Ridley et al., 2024) or transients that appear to transition between classifications over time (e.g., SN PTF11iqb, which observationally evolved from a Type IIn to Type IIP/L at late times; Smith et al., 2015).
@@
For this first release, we focus on ingesting and analyzing data from TDE candidates. We curate this dataset by directly scraping photometry from pre-existing publications, rather than re-reducing it ourselves. Observations of TDE candidates exist across the electromagnetic spectrum, making them a good starting point for the multiwavelength capabilities of OTTER. As a result, each transient currently in OTTER has, at least at one time, been classified as a TDE or TDE candidate. To illustrate the types of analyses facilitated by OTTER, we also present several worked examples focused on exploring the TDE population and comparing to theory.
[…]
This paper is organized as follows. In §II we describe the dataset schema, the data cleaning process, the database backend, and the classification flagging. In §III, we outline the website and the API, and describe how to contribute new datasets to the catalog as they get published. In §IV, we summarize the current state of the dataset. In §V, we outline a few applications of the OTTER dataset and API to the population of TDE candidates currently stored in OTTER. We conclude and discuss future directions in §VI. Appendix A lists the metadata and references for all of the transients currently in OTTER. Appendix B extends §II with a detailed description of the dataset schema. Appendix C includes additional plots of the current OTTER dataset.
II Methodology & Data Organization
@@
    A complete dataset of transient events that have at least one associated refereed publication, including their photometry and meta data (names, coordinates, distances, etc.). In this work we start with TDE candidate data but with the plan to expand OTTER to include other transient events in the future.
@@
    Multiple classification flags for each included transient, including an overall confidence flag and intuitive data filtering options, to make it easy for users to curate a sample optimal for their science case from the broader OTTER catalog.
  5. 5.
@@
    Cleaned Dataset: Similar to the OAC, and for longevity, we store all of our data in a human readable JSON format. The raw JSON files are currently available on GitHub [^2] https://github.com/astro-otter/otter-backups and will be made available on Zenodo as the dataset size grows. These JSON files are described in §II.1. To curate a clean dataset of TDE candidate photometry, we pull raw datasets from papers, public catalogs, and data uploaded by external users to the OTTER website and run them through our data cleaning pipeline, described in §II.3.
@@
    Backend Database: We use a database server to make the files programmatically accessible for both the API and the website. This is described in §II.2. This is hosted by Johns Hopkins’ SciServer  [^3] https://www.sciserver.org/.
@@
    Website: The OTTER website provides an easy view of the data available in OTTER. This is described in §III.1. This is also hosted by Johns Hopkins’ SciServer .
@@
    API: There is a Python API, built on the more general RESTful API [^4] https://docs.arangodb.com/stable/develop/http-api/, that allow user access to the cleaned dataset. This is described in §III.2 and in the API documentation [^5] https://astro-otter.readthedocs.io/. The API documentation also includes multiple Python Jupyter notebooks with examples of using the API to query the database [^6] https://astro-otter.readthedocs.io/en/latest/examples/welcome.html.
II.1 Data Schema
@@ II.2 Data Backend
II.3 Data Sources & Cleaning Pipeline
To create the initial OTTER TDE candidate catalog, we compile data from multiple existing catalogs and individual papers. In addition to cleaning the data into a standard format, each component of each dataset is appropriately cited with an ADS bibcode in the individual JSON files. All of these citations are given in Table LABEL:tab:refs. In §II.3.1 through §II.3.5 we describe dataset specific nuances for the data cleaning process.
@@ II.3 Data Sources & Cleaning Pipeline
Throughout the data cleaning process, we deduplicate identical data (metadata and photometry) by appending the new citation to the previously stored data. This ensures that all appropriate works are cited when the data is used. An intentional artifact of only deduplicating identical datasets is that it does not remove different reductions of the same photometry, allowing the user to choose which flux value they prefer. To prevent users from accidentally using multiple photometry points produced from different reductions of the same data, we provide a default deduplication algorithm as part of the Python API. See §III.2 for a discussion of this algorithm.
II.3.1 The Open TDE Catalog
The Open TDE Catalog was, like the Open Supernova Catalog, part of the AstroCats movement to create a standardized data format for public time domain data, including published and unpublished datasets (Guillochon et al., 2017; Auchettl et al., 2017). Until $\sim 2022$, this was the most comprehensive catalog of TDE data, with data on $\sim 100$ TDE candidates. The data organization in OTTER is influenced by the AstroCats movement, which also used the JSON format for their backend data storage.
@@ II.3.1 The Open TDE Catalog
II.3.2 Ultraviolet, Optical, and Infrared Data (UVOIR)
The so-called “Curated Optical TDE Catalog” is a continuously updated catalog of optically-selected TDE photometry available at https://github.com/sjoertvv/manyTDE, with the initial catalog and data curation methods described in Mummery et al. (2024b). This dataset includes more recent optical/UV/IR photometry for optically-selected TDEs, most of which were discovered after the deprecation of The Open TDE Catalog. Most of this photometry is from the Zwicky Transient Facility (ZTF; van Velzen et al., 2019) forced photometry server, WISE (Wright et al., 2010; Mainzer et al., 2011), and the Neil Gehrels Swift Ultraviolet/Optical Telescope (Swift UVOT), with some additional photometry from other optical all sky surveys (Shappee et al., 2014; Tonry et al., 2018c; Blagorodnova et al., 2017a; Gezari et al., 2012a). The catalog was initially assembled for the analysis performed in Mummery et al. (2024b), using methods outlined in van Velzen et al. (2019) and van Velzen et al. (2019), but has grown since then. All of the data is host subtracted and we store the appropriate host subtraction in the OTTER schema. Every time our data cleaning pipeline runs it pulls changes from The Curated Optical TDE Catalog and incorporates the photometry into OTTER.
@@ II.3.2 Ultraviolet, Optical, and Infrared Data (UVOIR)
Masterson et al. (2024) performed an archival search of WISE Infrared (IR) data for flares associated with potential tidal disruption events and found 12 new IR bright TDE candidates. We include their entire photometric dataset, spanning from IR to X-rays, in OTTER as well as specific metadata information like coordinates, redshift, and best-fit disruption time. This dataset was provided through private communication with the authors of Masterson et al. (2024).
@@ X VLA (8, 12]
II.3.5 The Transient Name Server
@@
II.4 Classification Confidence Flagging
Accurately and efficiently classifying transients, particularly rare transients like TDEs, remains a complex problem and a matter of intense discussion. For example, in the case of TDEs, while the gold standard in the field is often taken to be the existence of a high signal-to-noise optical spectrum with characteristic emission lines (e.g., Arcavi et al. 2014), this excludes TDE candidates that are discovered at other wavelengths (e.g., X-ray, radio) and do not show any optical emission — even though many of these objects are also widely accepted as TDEs (e.g., Auchettl et al., 2017; Anderson et al., 2020). To help with this ambiguity and facilitate the curation of “reliable” population sub-samples for different science cases, we provide numerous flags for the classifications associated with any given transient. The classification flags described in this subsection are applicable to most transients that will eventually be included in OTTER, including both TDEs and SNe. However, with the addition of some transient events (e.g., GRBs) in the future, other, more relevant, classification flags may become necessary. These flags also prepare us for a future where many transients will not be optically spectroscopically classified, in which case the classification may remain ambiguous indefinitely.
@@ II.4 Classification Confidence Flagging
Based on this flagging schema, we can outline the automated algorithm for deriving meta-level flags describing how each transient was classified. Figure 2 is a flow chart that shows this algorithm. Currently, we store three such meta-level flags. The first is simply the discovery method of the transient (discovery_method). The discovery_method flag is currently either “xray”, “uv”, “optical”, “ir”, or “radio”, but it could be extended to also include gamma-rays, neutrinos, or gravitational waves as future discovery methods. This flag can be particularly useful for selecting TDE samples as the discovery wavelength regime is tied to some astronomers’ confidence that the event is a TDE rather than, for example, a SNe.
@@
Figure 3 shows the distribution of the maximum confidence flags associated with each transient event (i.e., TDE or TDE candidate) stored in OTTER. This distribution demonstrates that most ($\sim 150$) the events in OTTER are classified from optical spectroscopy and about a third as many are only photometrically classified. The most pure sample of TDEs in OTTER probably has the flags unambiguous $\rm true$} and spec_classed $\geq 1$. However, selecting on these flags does bias the sample to optically-selected TDE candidates, even though there are other highly probable TDE candidates in the catalog. As a result, we leave the exact sample selection to the discretion of the individual user.
@@
III Data Availability & Contributing Data
As described in §II, two of the primary goals of OTTER are 1) allow for the data to be easily accessed and 2) make it easy to contribute data after it is published, with the hope that others will make this contribution part of their typical publishing workflow. As such, we developed two ways to interface with the OTTER dataset, neither of which require any account registration or login: 1) an open access website for easily viewing and uploading datasets and 2) an API for programmatic access to the data. In the following subsections, we describe this infrastructure.
@@ III.1 OTTER Website
Each individual transient page includes the names assigned to the transient, a table of its basic properties, and plots of its light curve and spectral energy distribution. A sample of one of these pages is shown in Figure 5. The properties table lists what we have deemed to be the “default” values for the various transient properties that are available. For most properties, if multiple values are reported in the literature, then the default value is determined by the value that has the most citations. There are two exceptions: 1) the default name of the transient, for which we select the International Astronomical Union convention (i.e. TNS or TNS-like) names over other transient aliases used in the literature (which are also listed in the table) and 2) the default classification, which is based on the highest confidence flag161616see §II.4 for a discussion of the confidence flagging and the biases that this default classification algorithm may introduce.. The light curve and spectral energy distribution plots are shown if photometry is publicly available.
III.1.1 Uploading Data via the OTTER Website